# Fast Penalized Generalized Estimating Equations for Large Longitudinal Functional Datasets

### v1

@@ C.1 Additional Simulations: Gaussian AR1
Table 12: Functional Coefficient Estimation Performance relative to initial FoSR fit, which assumes working independence correlation structure ($\text{RMSE}_{AR1}/\text{RMSE}_{FoSR}$). Cells contain the average of 300 replicates $\pm$ SEM (SEM$=0.00$ indicates a value $<0.01$).
$N$ $n_{i}$ One-Step GLS-AR1 GLS-Ind FoSR
0.25 0.5 0.75 0.25 0.5 0.75 0.25 0.5 0.75 0.25 0.5 0.75
25 5 $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$
25 $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$
100 $0.95$$\pm$   $0.00$ $0.95$$\pm$   $0.00$ $0.94$$\pm$   $0.00$ $0.95$$\pm$   $0.00$ $0.95$$\pm$   $0.00$ $0.94$$\pm$   $0.00$ $0.95$$\pm$   $0.00$ $0.96$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$
50 5 $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$
25 $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$
100 $0.94$$\pm$   $0.00$ $0.94$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.94$$\pm$   $0.00$ $0.94$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.94$$\pm$   $0.00$ $0.96$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $0.98$$\pm$   $0.00$
100 5 $0.98$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$
25 $0.96$$\pm$   $0.00$ $0.96$$\pm$   $0.00$ $0.96$$\pm$   $0.00$ $0.96$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.96$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.97$$\pm$   $0.00$ $0.98$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$ $1.00$$\pm$   $0.00$
100 $0.91$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.94$$\pm$   $0.00$ $0.95$$\pm$   $0.00$ $0.93$$\pm$   $0.00$ $0.96$$\pm$   $0.00$ $0.97$$\pm$   $0.00$
Table 13: Functional Coefficient Joint 95% Confidence Interval Coverage. Data are simulated with an AR1 correlation structure. Cells contain the average of 300 replicates $\pm$ SEM (SEM$=0.00$ indicates a value $<0.01$).
$N$ $n_{i}$ One-Step GLS-AR1 GLS-Ind FoSR
0.25 0.5 0.75 0.25 0.5 0.75 0.25 0.5 0.75 0.25 0.5 0.75
25 5 $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$
25 $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.88$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.88$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$
100 $0.87$$\pm$   $0.00$ $0.87$$\pm$   $0.00$ $0.85$$\pm$   $0.00$ $0.86$$\pm$   $0.00$ $0.87$$\pm$   $0.00$ $0.85$$\pm$   $0.00$ $0.87$$\pm$   $0.00$ $0.88$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.96$$\pm$   $0.00$ $0.96$$\pm$   $0.00$ $0.95$$\pm$   $0.00$
50 5 $0.91$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$
25 $0.90$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$
100 $0.86$$\pm$   $0.00$ $0.86$$\pm$   $0.00$ $0.84$$\pm$   $0.00$ $0.86$$\pm$   $0.00$ $0.86$$\pm$   $0.00$ $0.84$$\pm$   $0.00$ $0.86$$\pm$   $0.00$ $0.88$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.93$$\pm$   $0.00$ $0.94$$\pm$   $0.00$ $0.95$$\pm$   $0.00$
100 5 $0.91$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.92$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$
25 $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.88$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.88$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.90$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$ $0.99$$\pm$   $0.00$
100 $0.83$$\pm$   $0.00$ $0.84$$\pm$   $0.00$ $0.82$$\pm$   $0.00$ $0.84$$\pm$   $0.00$ $0.85$$\pm$   $0.00$ $0.83$$\pm$   $0.00$ $0.84$$\pm$   $0.00$ $0.86$$\pm$   $0.00$ $0.87$$\pm$   $0.00$ $0.89$$\pm$   $0.00$ $0.91$$\pm$   $0.00$ $0.94$$\pm$   $0.00$
Table 14: Functional Coefficient Pointwise 95% Confidence Interval Coverage. Data are simulated with an AR1 correlation structure. Cells contain the average of 300 replicates $\pm$ SEM (SEM$=0.00$ indicates a value $<0.01$).
$N$ $n_{i}$ One-Step GLS-AR1 GLS-Ind FoSR
0.25 0.5 0.75 0.25 0.5 0.75 0.25 0.5 0.75 0.25 0.5 0.75
25 5 $\phantom{0}16.54$$\pm$   $0.10$ $\phantom{0}16.21$$\pm$   $0.07$ $\phantom{0}24.79$$\pm$   $0.18$ $\phantom{00}8.35$$\pm$   $0.04$ $\phantom{00}8.20$$\pm$   $0.02$ $\phantom{0}12.08$$\pm$   $0.01$ $\phantom{00}9.69$$\pm$   $0.05$ $\phantom{00}9.55$$\pm$   $0.04$ $\phantom{0}12.26$$\pm$   $0.02$ $0.15$$\pm$   $0.00$ $0.15$$\pm$   $0.00$ $0.22$$\pm$   $0.00$
25 $\phantom{0}22.56$$\pm$   $0.15$ $\phantom{0}20.80$$\pm$   $0.12$ $\phantom{0}20.47$$\pm$   $0.11$ $\phantom{0}12.28$$\pm$   $0.10$ $\phantom{0}11.55$$\pm$   $0.08$ $\phantom{0}11.13$$\pm$   $0.05$ $\phantom{0}12.52$$\pm$   $0.11$ $\phantom{0}11.78$$\pm$   $0.09$ $\phantom{0}11.23$$\pm$   $0.07$ $0.28$$\pm$   $0.00$ $0.27$$\pm$   $0.00$ $0.26$$\pm$   $0.00$
100 $\phantom{0}45.85$$\pm$   $0.58$ $\phantom{0}39.32$$\pm$   $0.17$ $\phantom{0}41.68$$\pm$   $0.50$ $\phantom{0}34.17$$\pm$   $0.50$ $\phantom{0}29.34$$\pm$   $0.31$ $\phantom{0}28.54$$\pm$   $0.27$ $\phantom{0}30.27$$\pm$   $0.46$ $\phantom{0}26.13$$\pm$   $0.22$ $\phantom{0}26.24$$\pm$   $0.27$ $1.05$$\pm$   $0.02$ $0.72$$\pm$   $0.00$ $0.73$$\pm$   $0.01$
50 5 $\phantom{0}24.18$$\pm$   $0.15$ $\phantom{0}22.93$$\pm$   $0.14$ $\phantom{0}23.26$$\pm$   $0.06$ $\phantom{0}14.60$$\pm$   $0.09$ $\phantom{0}14.44$$\pm$   $0.08$ $\phantom{0}14.16$$\pm$   $0.03$ $\phantom{0}16.29$$\pm$   $0.09$ $\phantom{0}16.22$$\pm$   $0.10$ $\phantom{0}16.21$$\pm$   $0.08$ $0.22$$\pm$   $0.00$ $0.21$$\pm$   $0.00$ $0.21$$\pm$   $0.00$
25 $\phantom{0}33.37$$\pm$   $0.22$ $\phantom{0}32.74$$\pm$   $0.28$ $\phantom{0}33.42$$\pm$   $0.25$ $\phantom{0}21.95$$\pm$   $0.24$ $\phantom{0}21.54$$\pm$   $0.15$ $\phantom{0}22.68$$\pm$   $0.25$ $\phantom{0}21.62$$\pm$   $0.19$ $\phantom{0}21.32$$\pm$   $0.13$ $\phantom{0}22.30$$\pm$   $0.20$ $0.48$$\pm$   $0.00$ $0.48$$\pm$   $0.00$ $0.49$$\pm$   $0.00$
100 $\phantom{0}82.80$$\pm$   $1.03$ $\phantom{0}84.35$$\pm$   $1.11$ $\phantom{0}86.85$$\pm$   $1.19$ $\phantom{0}61.59$$\pm$   $0.72$ $\phantom{0}69.19$$\pm$   $1.03$ $\phantom{0}68.34$$\pm$   $1.02$ $\phantom{0}53.39$$\pm$   $0.65$ $\phantom{0}60.84$$\pm$   $0.92$ $\phantom{0}59.75$$\pm$   $0.90$ $1.54$$\pm$   $0.02$ $1.79$$\pm$   $0.03$ $1.66$$\pm$   $0.03$
100 5 $\phantom{0}37.37$$\pm$   $0.13$ $\phantom{0}36.62$$\pm$   $0.12$ $\phantom{0}37.55$$\pm$   $0.13$ $\phantom{0}26.49$$\pm$   $0.09$ $\phantom{0}26.45$$\pm$   $0.08$ $\phantom{0}26.23$$\pm$   $0.07$ $\phantom{0}27.42$$\pm$   $0.11$ $\phantom{0}28.07$$\pm$   $0.16$ $\phantom{0}27.67$$\pm$   $0.13$ $0.30$$\pm$   $0.00$ $0.31$$\pm$   $0.00$ $0.30$$\pm$   $0.00$
25 $\phantom{0}62.10$$\pm$   $0.80$ $\phantom{0}70.89$$\pm$   $0.20$ $\phantom{0}57.45$$\pm$   $0.47$ $\phantom{0}46.39$$\pm$   $0.48$ $\phantom{0}59.69$$\pm$   $0.61$ $\phantom{0}46.66$$\pm$   $0.57$ $\phantom{0}42.84$$\pm$   $0.30$ $\phantom{0}52.79$$\pm$   $0.25$ $\phantom{0}41.42$$\pm$   $0.40$ $0.77$$\pm$   $0.00$ $1.24$$\pm$   $0.00$ $0.76$$\pm$   $0.00$
100 $187.81$$\pm$   $3.17$ $175.73$$\pm$   $2.32$ $176.43$$\pm$   $2.67$ $159.26$$\pm$   $2.62$ $161.64$$\pm$   $2.81$ $144.59$$\pm$   $2.33$ $131.43$$\pm$   $5.45$ $122.23$$\pm$   $1.69$ $114.57$$\pm$   $1.50$ $3.17$$\pm$   $0.08$ $4.16$$\pm$   $0.09$ $3.09$$\pm$   $0.04$
Table 15: Fit Time (seconds) for the entire One-Step estimation procedure including 1) estimating the initial $\widehat{\bm{\beta}}_{\Lambda_{0}}^{(0)}$, 2) estimating correlation parameters, 3) tuning $\Lambda_{1}$, 4) calculating the One-Step $\widehat{\bm{\beta}}_{\Lambda_{1}}^{(1)}$, and 5) estimating $\widehat{\text{Var}}(\widehat{\bm{\beta}}_{\Lambda_{1}}^{(1)})$ and constructing CIs. Data are simulated with an AR1 correlation structure. Cells contain the average of 300 replicates $\pm$ SEM (SEM$=0.00$ indicates a value $<0.01$).
C.2 Gaussian Exchangeable
[…]
Table 16 shows the estimation performance and illustrates that the One-Step performs comparably to the GLS-Ex, suggesting the One-Step performs comparably to a fully-iterated GEE. Moreover the One-Step performs comparably to the marginal approach (Li et al.,, 2022) that models correlation across the functional domain (both within- and across longitudinal observations of the functional outcome). Thus, in these simulations, modeling correlation at each point $s$, is enough to capture efficiency gains. Table 18 shows that the pointwise coverage of the One-Step is often close to the nominal levels. In contrast, the Marginal approach is often highly anti-conservative for large $n_{i}$, a quality acknowledged in (Li et al.,, 2022). The GLS approaches are often highly conservative, while the FoSR is slightly more anti-conservative compared to the One-Step in smaller samples. Table 17 shows that the joint coverage tends to be slightly more conservative than the FoSR, but both are substantially less conservative than the GLS approaches.
$N$ $n_{i}$ One-Step GLS-Ex GLS-Ind Marginal
[…]

### v2

@@ C.3.2 Simulation Results
Table S11: Functional Coefficient Estimation Performance (RMSE) of each method relative to the pffr fit ($\text{RMSE}/\text{RMSE}_{\text{pffr}}$). Cells contain the average of 300 replicates $\pm$ SE (SE=0.00=0.00 indicates a value <0.01<0.01). The “1-Step” indicates a one-step was used for tuning and final coefficient estimation, and “Full-1step” indicates one-step tuning and a fully-iterated fGEE for final coefficient estimation, with the indicated working correlation. We indicate out-of-memory (35Gb) with symbol — .
Exchangeable
Method Gaussian
$N=25,\;n_{i}=5$
Li et. al 0.97 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.00 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.96 $\pm$ 0.00
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.03 $\pm$ 1 0.01
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 4 0.00
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.01 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.96 $\pm$ 0.00
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.04 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 4 0.00
$N=25,\;n_{i}=100$
Li et. al 1.05 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.02 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.99 $\pm$ 1 0.00
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.03 $\pm$ 0.01
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 3 0.00
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.02 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.99 $\pm$ 1 0.00
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.03 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 3 0.00
$N=50,\;n_{i}=5$
Li et. al 0.96 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.99 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.95 $\pm$ 0.01
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.02 $\pm$ 0.01
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 2 0.00
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.99 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.95 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.02 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 2 0.00
$N=50,\;n_{i}=100$
Li et. al 1.01 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.00 $\pm$ 1 0.00
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.99 $\pm$ 0.00
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.01 $\pm$ 0.00
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 1 0.00
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.00 $\pm$ 1 0.00
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.99 $\pm$ 0.00
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.01 $\pm$ 0.00
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 1 0.00
$N=500,\;n_{i}=5$
Li et. al 0.94 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.96 $\pm$ 0.00
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.95 $\pm$ 0.00
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.00 $\pm$ 1 0.00
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 0.99 $\pm$ 0.00
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.96 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.95 $\pm$ 0.00
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.00 $\pm$ 1 0.00
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 0.99 $\pm$ 0.00
$N=500,\;n_{i}=100$
Li et. al       —
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.00 $\pm$ 0.00
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 1.00 $\pm$ 0.01
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.00 $\pm$ 0.00
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 0.00
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.00 $\pm$ 0.00
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 1.00 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 1.00 $\pm$ 0.00
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 1.00 $\pm$ 0.00
Table S12: Pointwise 95% CI coverage from 300 replicates $\pm$ SE (SE=0.00=0.00 indicates a value <0.01<0.01). Table columns indicate if $\mathbf{R}^{*}_{\text{Lon}}$ had exchangeable or AR1 correlation. pffr (Wild) indicates CIs constructed for pffr coefficient estimates with quantiles obtained from a wild cluster bootstrap. pffr ($z_{1-\alpha/2}$) are standard Wald CIs constructed with Gaussian quantiles. We indicate out-of-memory (30Gb) with symbol — .
Exchangeable
Method Gaussian
$N=25,\;n_{i}=5$
Li et. al 0.95 $\pm$ 1 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.88 $\pm$ 0.02
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.92 $\pm$ 0.02
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.87 $\pm$ 0.02
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 0.91 $\pm$ 0.02
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.88 $\pm$ 0.02
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.92 $\pm$ 0.02
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.87 $\pm$ 0.02
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 0.91 $\pm$ 0.02
pffr (Wild) 0.97 $\pm$ 1 0.01
pffr (Sandwich) 0.88 $\pm$ 0.02
$N=25,\;n_{i}=100$
Li et. al 0.82 $\pm$ 0.02
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.94 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.94 $\pm$ 1 0.01
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 0.02
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 0.93 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.94 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.94 $\pm$ 1 0.01
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 0.02
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 0.93 $\pm$ 0.01
pffr (Wild) 0.97 $\pm$ 0.01
pffr (Sandwich) 0.92 $\pm$ 2 0.02
$N=50,\;n_{i}=5$
Li et. al 0.95 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 0.02
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.94 $\pm$ 0.01
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 1 0.01
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 0.94 $\pm$ 1 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 0.02
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.94 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 1 0.01
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 0.94 $\pm$ 1 0.01
pffr (Wild) 0.98 $\pm$ 1 0.01
pffr (Sandwich) 0.92 $\pm$ 1 0.02
$N=50,\;n_{i}=100$
Li et. al 0.80 $\pm$ 0.02
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.93 $\pm$ 0.01
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 0.01
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 0.94 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.93 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.93 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 0.94 $\pm$ 0.01
pffr (Wild) 0.98 $\pm$ 0.01
pffr (Sandwich) 0.92 $\pm$ 0.02
$N=500,\;n_{i}=5$
Li et. al 0.96 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.95 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.95 $\pm$ 0.01
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.94 $\pm$ 0.01
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 0.95 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.95 $\pm$ 0.01
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.95 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.94 $\pm$ 0.01
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 0.95 $\pm$ 0.01
pffr (Wild) 0.99 $\pm$ 0.01
pffr (Sandwich) 0.91 $\pm$ 0.02
$N=500,\;n_{i}=100$
Li et. al       —
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.95 $\pm$ 0.04
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.95 $\pm$ 0.04
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.96 $\pm$ 0.04
$\text{1-Step }I_{n_{i}}\otimes I_{L}$ 0.95 $\pm$ 0.04
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.95 $\pm$ 0.04
$\text{Full-1Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.95 $\pm$ 0.04
$\text{Full-1Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.96 $\pm$ 0.04
$\text{Full-1Step }I_{n_{i}}\otimes I_{L}$ 0.95 $\pm$ 0.04
pffr (Wild) 1.00 $\pm$ 0.00
pffr (Sandwich) 0.94 $\pm$ 0.04
Table S13: Relative pointwise CI width (mean $\pm$ SE) vs. pffr (Wild). We denote $UB^{(r)}(s)$ and $LB^{(r)}(s)$ and $UB_{\text{pffr}}^{(r)}(s)/LB_{\text{pffr}}^{(r)}(s)$ as the upper/lower bounds of the CIs (at $s$) of the indicated method and pffr, respectively. Below we report the average ratio $\frac{1}{(q+1)|\mathcal{S}|}\sum_{r=0}^{q}\sum_{s\in\mathcal{S}}\frac{UB^{(r)}(s)-LB(s)^{(r)}}{UB^{(r)}_{\text{pffr}}(s)-LB^{(r)}_{\text{pffr}}(s)}$ across 300 simulation replicates. Values <1<1 indicate narrower 95% CIs. Values with * indicate extreme outliers (from poor estimates) were removed from the average of that cell to avoid skewing results. We indicate out-of-memory (35Gb) with symbol — .
Exchangeable
Method Gaussian
$N=25,\;n_{i}=5$
Li et. al 0.69 $\pm$ 0.01
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.57 $\pm$ 0.00
$\text{1-Step }{\mathbf{R}}_{\text{Lon}}\otimes I_{L}$ 0.60 $\pm$ 0.00
$\text{1-Step }I_{n_{i}}\otimes{\mathbf{R}}_{\text{Fun}}$ 0.62 $\pm$ 0.00
[…]